# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

### v1

@@ Abstract
We propose a Frank–Wolfe (FW) algorithm with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions. This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property. Compared to existing FW algorithms, our assumptions are less restrictive. We establish convergence guarantees in various settings, such as sublinear to linear convergence rates, depending on the assumptions for convex and nonconvex objective functions. Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap. We also propose a variant of the away-step FW algorithm using Bregman distances. We establish global accelerated (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition over polytopes. Numerical experiments on nonnegative linear inverse problems, $\ell_{p}$ loss problems, phase retrieval, low-rank minimization, and nonnegative matrix factorization demonstrate that our proposed FW algorithms outperform existing methods.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a continuously differentiable function and $P\subset\mathbb{R}^{n}$ is a compact convex set. We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$. Usually, Problem (1.1) is solved by a variant of the projected gradient method (see, for example, [Nesterov2018-gj, Nocedal2006-sb]). However, even if $P$ is convex, it is not always possible to access the projection operator of $P$ or compute the projection efficiently. While interior point methods are a potential alternative approach to constrained optimization problems, their updates often require second-order information, which can be computationally expensive for large-scale problems. To alleviate the situation, Frank and Wolfe [Frank1956-cq] proposed the Frank–Wolfe (FW) algorithm (also known as the conditional gradient method), which is a projection-free first-order method. Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{argmin}_{v\in P}\langle a,v\rangle$. Linear minimization oracles are often much cheaper than projection oracles, as shown in [CP2021] (see also summary in [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods when the projection operation is not trivial. Additionally, FW algorithms tend to be numerically quite robust and stable due to the affine-invariance of the algorithm and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
The original FW algorithm is given in Algorithm 1.
@@
Input: Initial point $x_{0}\in P$, objective function $f$, step-size strategy $\gamma_{t}\in[0,1]$
@@ 4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
Related Work
In this section, we briefly review the work most closely related to this paper. The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably. Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte  [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax]. Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error. Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar complexity lower bound.
The classical FW algorithm is known to zigzag when approaching the optimal face containing the optimal solution $x^{*}$. This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterate sequence. Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm. Their analysis introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region. Similar analyses based on this constant have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021]. Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis.
Pedregosa et al. [Pedregosa2020-ay] introduced a powerful adaptive step-size strategy that dynamically approximates the local Lipschitz continuity of the function $f$. They proved that this strategy is at least as effective as the traditional short-step strategy, and it was later enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability. For nonconvex functions, Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, and Maskan et al. [Maskan2025-oa] incorporated the difference of convex functions (DC) into FW algorithms. For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al. [Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
Establishing the convergence rate of FW algorithms typically requires that $\nabla f$ is Lipschitz continuous, although exceptions exist, such as in the case of generalized self-concordant functions [CBP2021, SelfConcordantFW_2020]. Some first-order methods using the Bregman distance do not require the global Lipschitz continuity of $\nabla f$. Bolte et al. [Bolte2018-zt] proposed the Bregman proximal gradient algorithm and established its global convergence under the smooth adaptable property (see Definition 2.6). Hanzely et al. [Hanzely2021-sz] introduced accelerated Bregman proximal gradient algorithms. Takahashi et al. [Takahashi2022-ml] developed a Bregman proximal gradient algorithm for the DC optimization and its accelerated version. Subproblems of Bregman proximal gradient algorithms are often not solvable in closed form. Takahashi and Takeda [Takahashi2024-wz] used approximations to Bregman distances to solve subproblems in closed form. Rebegoldi et al. [Rebegoldi2018-pa] proposed an inexact version of the Bregman proximal gradient algorithm.
Contribution
Our contributions can be roughly summarized as follows:
Frank–Wolfe for Smooth Adaptable functions
We study Frank–Wolfe algorithms for smooth adaptable functions (also called: relatively smooth) functions. Some functions whose gradients are not Lipschitz continuous satisfy the smooth adaptable property with kernel generating distances (see Definition 2.1). The class of smooth adaptable functions appears in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt] phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh]. While the gradient of any $\mathcal{C}^{2}$ function over compact sets is Lipschitz continuous, an $\ell_{p}$ loss function is not $\mathcal{C}^{2}$ when $21<p<$ 2, i.e., its gradient is not Lipschitz continuous over compact sets. Moreover, functions in these applications are weakly convex on compact sets. We establish sublinear to linear convergence rates for the respective variants, which are compatible and in line with existing results.
Adaptive Bregman Step-Size Strategy
We propose an adaptive Bregman step-size strategy and establish its convergence results for convex and nonconvex optimization. Table 1 summarizes our contributions for the various cases that we consider. While Vyguzov and Stonyakin [Vyguzov2024-tp] also proposed an adaptive FW algorithm with the Bregman distance similar to ours for convex optimization problems. They established a linear convergence rate assuming that the objective function is relatively strongly convex, the optimal solution belongs to $\operatorname{int}P$, and the angle condition holds; in contrast, we establish linear convergence under much weaker assumptions. Beyond that, we also establish sublinear convergence for convex optimization for the case of general relatively smooth functions. Assuming that $f$ is convex, satisfies the Hölder error bound condition, and the optimal solution $x^{*}\in\operatorname{int}P$, we also provide the accelerated (up to linear) convergence, however, in a different way than [Vyguzov2024-tp] with fewer assumptions. For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence for our step-size strategy. These results include a special case that $\nabla f$ is Lipschitz continuous, i.e., $L$-smoothness. Without weak convexity, we prove the global convergence rate of the FW algorithm to a stationary point of (1.1), similar to [Lacoste-Julien2016-mj].
Away-Step FW Algorithm
We propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the Hölder error bound condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
Numerical Experiments
In numerical experiments, we have applied the FW algorithm with the adaptive Bregman step-size strategy to nonnegative linear systems [Bauschke2017-hg], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Patterson1934-sg, Patterson1944-nm], low-rank minimization [Dragomir2021-rv], and NMF [Gillis2020-dh]. Bauschke et al. [Bauschke2017-hg] showed applications to linear inverse problems to solve nonnegative linear systems. Maddison et al. [Maddison2021-nt] considered $\ell_{p}$ loss problems, and Kyng et al. [Kyng2015-uo] showed its application as Lipschitz learning on graphs. Bolte et al. [Bolte2018-zt] and Takahashi et al. [Takahashi2022-ml] applied Bregman proximal algorithms to phase retrieval, Dragomir et al. [Dragomir2021-rv] applied them to low-rank minimization, and Mukkamala and Ochs [Mukkamala2019-mk] and Takahashi et al. [Takahashi2024-cu] applied them to NMF. Throughout these applications, we compare our proposed algorithms with existing FW algorithms and the mirror descent algorithm [Nemirovskij1983-sm] as well as away-step FW algorithms (see e.g. [Braun2022-ew, GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax]).
Table 1: Convergence rates for the Frank–Wolfe algorithm. For nonconvex functions, convergence is measured using the Frank-Wolfe gap, $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap, $f(x_{t})-f^{*}\leq\epsilon$. Weak convexity can be replaced by twice continuous differentiability (see Proposition 2.10). Convergence rates for the weakly convex optimization locally hold. AFW, HEB, and quad. denote the away-step Frank–Wolfe algorithm, the Hölder error bound condition with $q\geq 1$, and the quadratic growth condition, respectively.
FW Assumptions Convergence rate
$f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu\in(0,1]$)
any convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2024-tp], Thm. 4.2)

### v3

# Fast Frank–Wolfe Algorithms with Adaptive Bregman Step-Size for Weakly Convex Functions

@@ Abstract
We propose a Frank–Wolfe (FW) algorithm with an adaptive Bregman step-size strategy for smooth adaptable (also called: relatively smooth) (weakly-) convex functions. This means that the gradient of the objective function is not necessarily Lipschitz continuous, and we only require the smooth adaptable property. Compared to existing FW algorithms, our assumptions are less restrictive. We establish convergence guarantees in various settings, such as sublinear to linear convergence rates, depending on the assumptions for convex and nonconvex objective functions. Assuming that the objective function is weakly convex and satisfies the local quadratic growth condition, we provide both local sublinear and local linear convergence regarding the primal gap. We also propose a variant of the away-step FW algorithm using Bregman distances over polytopes. We establish global faster (up to linear) convergence for convex optimization under the Hölder error bound condition and its local linear convergence for nonconvex optimization under the local quadratic growth condition. Numerical experiments demonstrate that our proposed FW algorithms outperform existing methods.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a continuously differentiable function and $P\subset\mathbb{R}^{n}$ is a compact convex set. We are interested in both convex and nonconvex $f$ and assume that we have first-order oracle access to $f$, i.e., given $x\in\mathbb{R}^{n}$, we can compute $\nabla f(x)$. The Frank–Wolfe (FW) algorithm is a projection-free first-order method. Instead of requiring access to a projection oracle, the FW algorithm requires only access to a so-called linear minimization oracle (LMO), which for a given linear function $a\in\mathbb{R}^{n}$ computes $y\in\operatorname*{argmin}_{v\in P}\langle a,v\rangle$. LMOs are often much cheaper than projection oracles, as shown in [CP2021] (see also [Braun2022-ew, Table 1.1]), so that, in practice, FW algorithms are often faster than projected gradient methods even when the projection operation is not trivial. Additionally, FW algorithms tend to be numerically quite robust and stable due to their affine-invariance and can be also used, e.g., to provide theoretical guarantees for the approximate Carathéodory problem [cp2019approxCara].
Overview of Existing FW Studies:
The FW algorithm was originally proposed by Frank and Wolfe [Frank1956-cq] and was independently rediscovered and extended by Levitin and Polyak [Levitin1966-sm] as the conditional gradient method; we will use these terms interchangeably. Canon and Cullum [Canon1968-sr] established an initial lower bound for the convergence rate of the FW algorithm, which was later improved by GuéLat and Marcotte [GueLat1986-bg], which also provided the first analysis of the away-step FW algorithm by Wolfe [Wolfe1970-ax]. Jaggi [Jaggi2013-oz] provided a more detailed convergence analysis of the FW algorithm, establishing a new lower bound that demonstrated a trade-off between sparsity and error. Concurrently, Lan [lan2013complexity] examined the complexity of linear programming-based first-order methods, establishing a similar lower bound. For a comprehensive review of FW algorithms, we refer the interested reader to the survey by Braun et al. [Braun2022-ew] and the brief introduction by Pokutta [Pokutta2024-qu].
The classical FW algorithm (Algorithm 1) has been improved in two main directions to enhance both performance and convergence. One focuses on refining the step-size rule. The short-step strategy computes $\gamma_{t}$ using the Lipschitz constant $L$ of $\nabla f$. Pedregosa et al. [Pedregosa2020-ay] introduced an adaptive step-size strategy that dynamically estimates $L$. They proved that this strategy is at least as effective as the short-step strategy. It was enhanced by Pokutta [Pokutta2024-qu] to improve numerical stability.
Another direction is to modify the classical FW algorithm to eliminate the zigzag behavior when approaching the optimal face containing the optimal solution $x^{*}$. This behavior inspired Wolfe [Wolfe1970-ax] to propose the away-step FW algorithm in 1970 that shortcuts the zigzagging by removing atoms that slow down the iterative sequence. Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] showed the linear convergence of the away-step FW algorithm. Their analysis***Similar analyses based on the pyramidal width have been developed for other advanced variants, such as the pairwise FW algorithm [Lacoste-Julien2015-gb], Wolfe’s algorithm [Lacoste-Julien2015-gb], the blended conditional gradient method [BPTW2018], and blended pairwise conditional gradient methods and similar variants [CP2020boost, TTP2021]. Even Nesterov-style acceleration is possible under weak assumptions [CDP2019] building upon this analysis. introduced a geometric constant, the pyramidal width, that measures the conditioning of the polytope $P$, representing the feasible region.
Our Research Idea:
To derive the convergence rate of the above-mentioned FW algorithms, previous studies have typically required $L$-smoothness, i.e., that $\nabla f$ is Lipschitz continuous, and the convexity for the objective function $f$, although exceptions exist.Quite recently, Vyguzov and Stonyakin [Vyguzov2025-ha] proposed a FW algorithm with the Bregman distance, similar to ours, yet it is restricted to convex problems and provides only partial convergence guarantees. These assumptions often narrow the application of FW algorithms. Even if they do not satisfy $L$-smoothness, there are many functions that satisfy the $L$-smooth adaptable property ($L$-smad) with kernel generating distances $\phi$ (see Definition 2.1) by choosing $\phi$ well to match the function $f$. Since the $L$-smad property is consistent with $L$-smoothness when $\phi=\frac{1}{2}\|\cdot\|^{2}$, the class of $L$-smad functions includes the $L$-smooth function class. Other first-order algorithms, such as the Bregman proximal gradient algorithm, are often analyzed under the $L$-smad property (see e.g., [Bolte2018-zt, Hanzely2021-sz, Rebegoldi2018-pa, Takahashi2022-ml, Takahashi2024-wz, Yang2025-ls]). $L$-smad functions appear in many applications, such as nonnegative linear inverse problems [Bauschke1997-vk, Takahashi2024-wz], $\ell_{p}$ loss problems [Kyng2015-uo, Maddison2021-nt], phase retrieval [Bolte2018-zt, Takahashi2022-ml], nonnegative matrix factorization (NMF) [Mukkamala2019-mk, Takahashi2024-cu], and blind deconvolution [Takahashi2023-uh]. Any $\mathcal{C}^{2}$ function is locally $L$-smooth over compact sets, but the resulting Lipschitz constant $L$ can be overly conservative, hindering practical performance. Functions outside $\mathcal{C}^{2}$ may not satisfy $L$-smoothness at all. Moreover, functions in the above-mentioned applications are weakly convex on compact sets. For nonconvex functions, although Lacoste-Julien [Lacoste-Julien2016-mj] established sublinear convergence, little else has been done, and simply tracing existing research does not achieve convergence rates better than sublinear convergence. Then the following question naturally arises:
Is it possible to loosen the $L$-smoothness and convexity assumptions often made in FW algorithm studies while still achieving the theoretical guarantee of linear convergence rates?
Contribution:
The answer to the above question is yes. We develop FW algorithms for $L$-smad $f$ and give theoretical guarantees of linear convergence under weaker assumptions (Hölder error bound (HEB) condition with parameter $q\geq 1$ for convex $f$ and the special case with $q=2$ called quadratic growth for nonconvex $f$) than the strong convexity assumption that is often made when linear convergence is shown for FW algorithms. The convexity of $f$ is also weaken to weakly convex, i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex with some $\rho>0$ (see, e.g.[Davis2018-bb, Examples 3.1 and 3.2] and [Nurminskii1975-vp]). Moreover, it is easy to see that any $\mathcal{C}^{2}$ function defined on a compact set is weakly convex [Vial1983-oo, Proposition 4.11] and [Wu2007-dt]; since $\rho$ is only needed for the theoretical guarantee and not for the algorithm, there is no need for its estimation, and considering weakly-convex $f$ greatly expands the range of applications.
Table 1 summarizes our contributions for the various cases that we consider. Relaxing the assumption from $L$-smoothness to the $L$-smad property makes the construction of a FW algorithm and its theoretical guarantee much more difficult. This is because we use the Bregman distance, which is an extension of the Euclidean distance, for the $L$-smad property. While (2.1) is simple with $\nu=1$ when using the Euclidean distance $\phi=\frac{1}{2}\|\cdot\|^{2}$, $\nu>0$ cannot be eliminated for the general Bregman distance; $1+\nu$ represents its scaling exponent. $\nu$ must be estimated to use in the algorithm since its exact value is unknown. Therefore, the extension to an $L$-smad $f$ is not straightforward.
  • We establish sublinear convergence for the case of convex $L$-smad functions. Assuming that $f$ is convex, satisfies the HEB condition, and the minimizer $x^{*}\in\operatorname{int}P$, we provide the faster (up to linear) convergence. In this setting, our algorithm always converges linearly if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$.As discussed in Sec. 2, linear convergence with the Euclidean distance for convex optimization is shown in [garber2015faster] under the quadratic growth over strongly convex sets. This corresponds to the $q=1+\nu$ case with $\nu=1$. For $q>1+\nu$, if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$, linear convergence holds; otherwise, $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$, which is faster than existing sublinear rates. A similar argument also applies to the nonconvex case ($q=2$) and the away-step FW algorithm.
  • For nonconvex optimization, assuming that $f$ is weakly convex and satisfies the local quadratic growth condition, we establish local linear convergence. To our knowledge, this is the first proposal of a FW algorithm with linear convergence guarantees for a certain type of nonconvex optimization problem, albeit of a limited type. These results are also new for the standard setting: $L$-smooth $f$. Without weak convexity, we prove a global sublinear convergence to a stationary point of (1.1) for nonconvex $L$-smad $f$.
  • We also propose a variant of the away-step FW algorithm with the Bregman distance and establish its linear convergence under the HEB condition for convex optimization and under the local quadratic growth condition for weakly convex optimization.
Table 1: Convergence rates for Frank–Wolfe algorithms. Our convergence rates are generalizations of existing rates. Moreover, our convergence rate is better than $\mathcal{O}(\epsilon^{-1})$ in some cases. For nonconvex functions, convergence is measured using the Frank–Wolfe gap $\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\leq\epsilon$, instead of the primal gap $f(x_{t})-f^{*}\leq\epsilon$. Convergence rates for the weakly convex optimization locally hold.
Assumptions Convergence rate
FW $f$ convexity $f$ growth $x^{*}\in\operatorname{int}P$ $P$ polytope $L$-smooth $L$-smad ($\nu>0$)
Alg.1 (any step-size) convex $\mathcal{O}(\epsilon^{-1})$ [Jaggi2013-oz] $\mathcal{O}(\epsilon^{-1/\nu})$ ([Vyguzov2025-ha], Thm.C.1)
Alg.1 (short & adapt.) convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$ [^2] It holds if the exponent of the HEB condition $q$ equals the scaling exponent of the Bregman distance $1+\nu$, i.e., $q=1+\nu$ ($(\nu,q)=(1,2)$ for the away-step FW algorithm). For $q>1+\nu$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(1+\nu-q)/\nu q})$.(Thm. 4.2)
away-step FW convex $q$-HEB $\mathcal{O}(\log\epsilon^{-1})$11footnotemark: 1[KDP2018jour] $\mathcal{O}(\log\epsilon^{-1})$22footnotemark: 2(Thm. 4.4)
Alg.1 (any step-size) weak 2-HEB $\mathcal{O}(\epsilon^{-1})$ $\mathcal{O}(\epsilon^{-1/\nu})$ (Thm. D.3)
Alg.1 (short & adapt.) weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$ [^4] It holds if $\nu=1$. For $\nu\in(0,1)$, it also holds if $t\leq t_{0}$ with some $t_{0}\in\mathbb{N}$; otherwise $\mathcal{O}(\epsilon^{(\nu-1)/2\nu})$.(Thm. 5.2)
away-step FW weak 2-HEB $\mathcal{O}(\log\epsilon^{-1})$ $\mathcal{O}(\log\epsilon^{-1})$44footnotemark: 4(Thm. 5.3)
Alg.1 (any step-size) $\mathcal{O}(\epsilon^{-2})$ [Lacoste-Julien2016-mj] $\mathcal{O}(\epsilon^{-1-1/\nu})$ (Thm. D.1)
2 Preliminaries
[…]
Let $C$ be a nonempty open convex subset of $\mathbb{R}^{n}$.
  1. (i)
    $\phi$ is proper, lower semicontinuous, and convex, with $\operatorname{dom}\phi\subset\operatorname{cl}C$ and $\operatorname{dom}\partial\phi=C$;
  2. (ii)
Definition 2.2 (Bregman Distance [Bregman1967-rf]).
Given $\phi\in\mathcal{G}(C)$, a Bregman distance $D_{\phi}:\operatorname{dom}\phi\times\operatorname{int}\operatorname{dom}\phi\to\mathbb{R}_{+}$ associated with $\phi$ is defined by $D_{\phi}(x,y)=\phi(x)-\phi(y)-\langle\nabla\phi(y),x-y\rangle\geq 0$.
We use $D_{f}$ in Definition 2.2 for nonconvex $f$ instead of convex $\phi$ for brevity. In the remainder of the paper, if not stated otherwise, we will assume that as such for a given strictly convex $\phi$ and $C$, there exists $\nu>0$ such that the following holds for all $x,y\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$ (see also Lemma A.3, which shows that a weaker one holds)):
Let $\phi\in\mathcal{G}(C)$ and let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C\equiv\operatorname{int}\operatorname{dom}\phi$. The pair of functions $(f,\phi)$ is said to be $L$-smooth adaptable (for short: $L$-smad) on $C$ if there exists $L>0$ such that $L\phi-f$ and $L\phi+f$ are convex on $C$.
Lemma 2.4 (Primal progress from the smooth adaptable property).
Let the pair of functions $(f,\phi)$ be $L$-smad with a strictly convex function $\phi\in\mathcal{G}(C)$ and let $x_{+}=(1-\gamma)x+\gamma v$ with $x,v\in\operatorname{int}\operatorname{dom}\phi$ and $\gamma\in[0,1]$. Then it holds:
$f(x)-f(x_{+})\geq\gamma\langle\nabla f(x),x-v\rangle-L\gamma^{1+\nu}D_{\phi}(v,x).$ (2.2)
Often the point $v\in P\subset\operatorname{int}\operatorname{dom}\phi$ in Lemma 2.4 is chosen as a Frank–Wolfe vertex, i.e., $v\in\operatorname*{argmax}_{u\in P}\langle\nabla f(x),x-u\rangle$, but other choices, e.g., those arising from away-directions from the away-step FW algorithm, are also possible, as we will see in Section 3.2.
In [garber2015faster] (see also [Garber2020-tr]) the linear convergence of the FW algorithm for convex optimization under the quadratic growth condition, which is a weaker assumption than assuming strong convexity, was established over strongly convex sets and later generalized to uniformly convex sets in [KAP2020] as well as to conditions weaker than quadratic growth in [KDP2018, KDP2018jour]. Local variants of these notions, as necessary, e.g., for the nonconvex case, have been studied in [KAP2021] in the context of FW algorithms. For $f:\mathbb{R}^{n}\to[-\infty,+\infty]$, let $[f\leq\zeta]:=\{x\in\mathbb{R}^{n}\,|\,f(x)\leq\zeta\}$ be a $\zeta$-sublevel set of $f$ for some $\zeta\in\mathbb{R}$.
Let $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ be a proper lower semicontinuous function and $P\subset\mathbb{R}^{n}$ be a compact convex set. Let $\mathcal{X}^{*}\neq\emptyset$ be the set of optimal solutions, i.e., $\mathcal{X}^{*}:=\operatorname*{argmin}_{x\in P}f(x)$, and let $f^{*}=\min_{x\in P}f(x)$ and $\zeta>0$. The function $f$ satisfies the Hölder error bound (HEB) condition on $P$ if there exists constants $q\geq 1$ and $\mu>0$ such that, for any $x\in P$,
The HEB condition shows sharpness bounds on the primal gap (see [bolte17, alex17sharp]), which has also been extensively analyzed for FW algorithms [KDP2018, KDP2018jour]. Convergence for nonconvex optimization under the quadratic growth or HEB condition was established in [Davis2024-rm, Davis2018-bb, Davis2024-og]. We employ the HEB condition for convex optimization, while we use the local quadratic growth for nonconvex optimization.
3 Proposed Bregman FW Algorithms
Throughout this paper, we make the following assumptions.
Assumption 3.1.
  1. (i)
    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$;
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is proper and lower semicontinuous with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$;
  3. (iii)
    The pair $(f,\phi)$ is $L$-smad on $P$;
  4. (iv)
    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz]. The strict convexity of $\phi$ ensures the existence of $\nu$ by Lemma A.3 and is satisfied by many kernel generating distances (see, e.g., [Bauschke2017-hg, Bauschke1997-vk], and [Dhillon2008-zz, Table 2.1]). Assumption 3.1(iv) is standard for FW algorithms, and $P\subset C$ is natural ensuring that $D_{\phi}$ is well-defined on $P$. The original FW algorithm is given in Algorithm 1.
@@ 3 Proposed Bregman FW Algorithms
Input: Initial point $x_{0}\in P$, objective function $f$, step-size $\gamma_{t}\in[0,1]$
@@ 4     $x_{t+1}\leftarrow(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$
In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and $f^{*}=f(x^{*})$. We recall a key property for analyzing convergence rates in convex optimization and derive its counterpart in weakly convex optimization.
Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap for convex $f$ [Pokutta2024-qu, Lemma 4.1]).
Suppose that Assumptions 3.1(ii), (iv), and that $f$ is convex. Then for all $x\in P$, it holds:
Lemma 3.3 (Primal gap, dual gap, and Frank–Wolfe gap for weakly convex $f$).
Suppose that Assumptions 3.1(ii), (iv), and that $f$ is $\rho$-weakly convex. Then for all $x\in P$, it holds:
$f(x)-f^{*}\leq\langle\nabla f(x),x-x^{*}\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}\leq\max_{v\in P}\langle\nabla f(x),x-v\rangle+\frac{\rho}{2}\|x-x^{*}\|^{2}.$ (3.2)
3.1 Bregman Step-Size Strategy for $\gamma_{t}$
Bregman Short Step-Size:
Assume that $(f,\phi)$ is $L$-smad. Using the optimality condition of the right-hand side of (2.2) in terms of $\gamma$, we find
$\gamma=\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x)-f(x_{+})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x),x-v\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v,x))^{1/\nu}}.$
Theoretically, when $x,v\in P$ and $\gamma\in[0,1]$, the Frank–Wolfe step $x_{+}=(1-\gamma)x+\gamma v\in P$ is well-defined. We update $\gamma=\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{L(1+\nu)D_{\phi}(v,x)}\right)^{1/\nu},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$). This step-size strategy provides the progress lemma.
Lemma 3.4 (Progress lemma from the Bregman short step-size).
Suppose that Assumption 3.1 holds. Define $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$. Then if $x_{t+1}\in\operatorname{dom}f$, $f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ for $0\leq\gamma_{t}\leq\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu}$.
Output : Estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$